# Disentangled and Self-Explainable Node Representation Learning

### v1

@@ 5 Conclusions
We present DiSeNE, a novel framework for generating self-explainable unsupervised node embeddings. To build our framework, we design new objective functions that ensure structural faithfulness, dimensional explainability, and structural disentanglement. Unlike traditional GNN explanation methods that typically extract a subgraph from a node’s local neighborhood, DiSeNE introduces a paradigm shift by learning node embeddings where each dimension captures an independent structural feature of the input graph. Additionally, we propose new metrics to evaluate the human interpretability of explanations, analyze the influence of spatial structures and node positions on latent features, and apply post-hoc feature attribution methods to derive task-specific instance-wise explanations.
@@ Acknowledgments
S. Piaggesi acknowledges support by the European Community program under the funding schemes: G.A. 834756 “XAI: Science and technology for the eXplanation of AI decision making”.
Appendix A Appendix / supplemental material
A.1 Dataset statistics
@@
Cora Wiki FB PPI Ring-Cl SBM BA-Cl ER-Cl
# nodes 2,485 2,357 4,039 3,480 320 320 640 640
# edges 5,069 11,592 88,234 53,377 1,619 1,957 3,138 4,196
# clusters 28 18 16 9 32 32 32 32
density 0.002 0.004 0.011 0.009 0.032 0.038 0.015 0.021
clust. coeff. 0.238 0.383 0.606 0.173 0.807 0.561 0.486 0.456
A.2 Training settings
@@
  • For DeepWalk [PARS14], we train Node2Vec [^5] https://github.com/eliorc/node2vec algorithm for 5 epochs with the following parameters: p$=1$, q$=1$, walk_length$=20$, num_walks$=10$, window_size$=5$.
@@
  • For InfWalk [^6] https://github.com/schariya/infwalk [CM20], a matrix factorization-based method linked to DeepWalk and spectral graph embeddings, we set the same value window_size$=5$ used for DeepWalk.
@@
  • In GraphAE [SHV21], we optimize a 1-layer GCN encoder with a random-walk loss setting analogous to DeepWalk. The model is trained for 50 iterations using Adam optimizer and learning rate of 0.01.
@@
  • in GraphSAGE [^7] https://github.com/pyg-team/pytorch_geometric/blob/master/examples/graph_sage_unsup.py [KO21], we optimize a 2-layer SAGE encoder with mean aggregation and with a random-walk loss setting analogous to DeepWalk. The model is trained for 50 iterations using Adam optimizer, learning rate of 0.01.
@@
  • Dine [^8] https://www.github.com/simonepiaggesi/dine [PKPA24], autoencoder-based post-processing process trained for $2000$ iterations, and learning rate of $0.1$. Input embeddings are from DeepWalk and GAE methods, tuning the input embedding size in the list $[8,16,32,64,128,256,512]$.
@@
  • DiSe-FCAE and DiSe-GAE trained for 50 iterations using Adam optimizer and learning rate of 0.01. Random walk sampling follows the same setting as DeepWalk, GraphAE and GraphSAGE.
A.3 Downstream Tasks
@@
We tested link prediction for the datasets reported in the main paper. For node classification, we tested PPI and other benchmark biological datasets in multi-label setting [ZDHK]: the PCG dataset for the protein phenotype prediction, the HumLoc, and EukLoc datasets for the human and eukaryote protein subcellular location prediction tasks, respectively. Characteristics of additional biological datasets are reported in Table A2. In Figure A1 we report AUC-PR scores for link prediction and node classification in real-world graph data. Generally, scores increase with the number of latent embedding dimensions. Tables A3 and A4 show the maximum scores for link prediction and node classification, demonstrating that our approach can consistently achieve reasonable performances.
@@
Figure A1: Downstream tasks results on real-world datasets with varying feature dimensions size.
@@
DeepWalk .892$\pm$.005 .927$\pm$.002 .990$\pm$.001 .794$\pm$.002
GraphAE .911$\pm$.003 .950$\pm$.001 .994$\pm$.001 .916$\pm$.001
InfWalk .923$\pm$.003 .936$\pm$.002 .941$\pm$.006 .854$\pm$.003
GraphSAGE .913$\pm$.005 .944$\pm$.002 .991$\pm$.001 .892$\pm$.003
DW+Dine .896$\pm$.004 .931$\pm$.003 .987$\pm$.001 .817$\pm$.004
GAE+Dine .926$\pm$.001 .957$\pm$.003 .992$\pm$.002 .919$\pm$.002
DiSe-FCAE .856$\pm$.007 .911$\pm$.004 .977$\pm$.001 .884$\pm$.002
DiSe-GAE .885$\pm$.002 .947$\pm$.002 .993$\pm$.006 .913$\pm$.001
@@
DeepWalk .476$\pm$.003 .210$\pm$.001 .314$\pm$.012 .241$\pm$.010
GraphAE .517$\pm$.003 .241$\pm$.001 .336$\pm$.004 .249$\pm$.005
InfWalk .442$\pm$.001 .207$\pm$.002 .287$\pm$.004 .212$\pm$.003
GraphSAGE .506$\pm$.001 .231$\pm$.002 .316$\pm$.004 .237$\pm$.011
DW+Dine .488$\pm$.002 .217$\pm$.001 .308$\pm$.004 .231$\pm$.008
GAE+Dine .526$\pm$.001 .241$\pm$.001 .333$\pm$.006 .234$\pm$.008
DiSe-FCAE .498$\pm$.001 .233$\pm$.003 .291$\pm$.006 .230$\pm$.006
DiSe-GAE .518$\pm$.001 .242$\pm$.004 .315$\pm$.003 .238$\pm$.006
A.4 Results for synthetic datasets
In Figure A2 we plot results for comprehensibility and sparsity, on the top and the bottom respectively, on synthetic datasets. Generally, DiSe-FCAE outperforms DiSe-GAE and the other competitors in all the datasets. In Figure A3 we plot results for overlap consistency and positional coherence, on the top and the bottom respectively, on synthetic datasets. For the overlap metric, DiSe-FCAE and DiSe-GAE consistently outperform the competitors, especially with more than 8 dimensions where they achieve almost perfect overlap. For the positional metric, the competitors GAE+Dine and DW+Dine slightly outperform DiSe methods, especially in large dimensions, while DeepWalk also show good results.
@@ A.4 Results for synthetic datasets
Figure A2: Comprehensibility and sparsity results on synthetic datasets with varying feature dimensions size.

### v2

@@ 5 Conclusions
We present DiSeNE, a novel framework for generating self-explainable unsupervised node embeddings. To build our framework, we design new objective functions that ensure connectivity preservation, dimensional explainability, and structural disentanglement. Unlike traditional GNN explanation methods that typically extract a subgraph from a node’s local neighborhood, DiSeNE introduces a paradigm shift by learning node embeddings where each dimension captures an independent structural feature of the input graph. Additionally, we propose new metrics to evaluate the human interpretability of explanations, analyze the influence of spatial structures and node positions on latent features, and apply post-hoc feature attribution methods to derive task-specific instance-wise explanations. Our results show that interpretable node representations for graphs can be obtained by disentangling topological substructures across embedding dimensions. Additionally, the most important node features identified by post-hoc techniques aligns with the true explanation subgraphs. These findings mark a significant step toward human-centric evaluations of node embeddings, pointing to critical directions for future work in advancing human-in-the-loop validations in graph feature learning.
@@ Acknowledgments
S. Piaggesi acknowledges support from the European Community program under the funding schemes: ERC-2018-ADG G.A. 834756 “XAI: Science and technology for the eXplanation of AI decision making”.
Appendix
@@
Cora Wiki FB PPI Ring-Cl SBM BA-Cl ER-Cl Tr-Cl Tr-Gr
# nodes 2,485 2,357 4,039 3,480 320 320 640 640 831 799
# edges 5,069 11,592 88,234 53,377 1,619 1,957 3,138 4,196 2,081 972
# clusters/motifs 28 18 16 9 32 32 32 32 32 32
density 0.002 0.004 0.011 0.009 0.032 0.038 0.015 0.021 0.006 0.003
clust. coeff. 0.238 0.383 0.606 0.173 0.807 0.561 0.486 0.456 0.360 0.002
Appendix A Training Settings
@@
For DeepWalk (Perozzi et al., 2014), we train Node2Vec [^6] https://github.com/eliorc/node2vec algorithm for 5 epochs with the following parameters: p=1, q=1, walk_length=20, num_walks=10, window_size=5.
@@
For InfWalk [^7] https://github.com/schariya/infwalk (Chanpuriya & Musco, 2020), a matrix factorization-based method linked to DeepWalk and spectral graph embeddings, we set the same value window_size=5 used for DeepWalk.
@@
In GraphAE (Salha et al., 2020), we optimize a 1-layer GCN encoder with a random-walk loss setting analogous to DeepWalk. The model is trained for 50 iterations using Adam optimizer and learning rate of 0.01.
@@
In GraphSAGE [^8] https://github.com/pyg-team/pytorch_geometric/blob/master/examples/graph_sage_unsup.py (Hamilton et al., 2017), we optimize a 2-layer SAGE encoder with mean aggregation and with a random-walk loss setting analogous to DeepWalk. The model is trained for 50 iterations using Adam optimizer, learning rate of 0.01.
In DGLFRM [^9] https://github.com/nikhil-dce/SBM-meet-GNN (Mehta et al., 2019), we optimize a 2-layer GCN encoder with the standard hyperparameters for variational inference: $\alpha_{0}$=10, $\lambda_{prior}$=0.5 and $\lambda_{post}$=1. The hidden embedding size is tuned in the list $[8,16,32,64,128,256,512]$. The model is trained for 300 iterations using Adam optimizer, learning rate of 0.01.
@@
Dine [^10] https://www.github.com/simonepiaggesi/dine (Piaggesi et al., 2024), autoencoder-based post-processing process trained for $2000$ iterations, and learning rate of $0.1$. Input embeddings are from DeepWalk and GAE methods, tuning the input embedding size in the list $[8,16,32,64,128,256,512]$.
@@
DiSe-FCAE and DiSe-GAE trained for 50 iterations using Adam optimizer and learning rate of 0.01. The hidden embedding size is tuned in the list $[8,16,32,64,128,256,512]$. Random walk sampling follows the same setting as DeepWalk, GraphAE and GraphSAGE.
GNNExplainer [^11] https://pytorch-geometric.readthedocs.io/en/latest/generated/torch_geometric.explain.algorithm.GNNExplainer.html#torch_geometric.explain.algorithm.GNNExplainer is trained for 30 epochs for each test node, while PGExplainer [^12] https://pytorch-geometric.readthedocs.io/en/latest/generated/torch_geometric.explain.algorithm.PGExplainer.html#torch_geometric.explain.algorithm.PGExplainer is trained for 5 epochs on trained nodes before being applied on test nodes. Moreover, since PGExplainer is based on edge masks, we derive node masks for that model with the average mask value from incident edges.
Graph explainers are applied on top of the following GNN models trained on node classification (two-layer for clique-based data and three-layer for grid-based data): GCN (Wu et al., 2019), GraphSAGE (Hamilton et al., 2017), and GATv2 (Brody et al., 2022). All the graph models (not the explainers) are tuned by searching the best output embedding size from the list $[2,4,8,16,32,64,128]$, as the input to the classification layer.
Appendix B Algorithm Complexity
Space and time complexity of DiseNE can be analyzed by looking at the pseudo-code in Algorithm 1. Part of the complexity depends on the complexity of the encoder. Here, we assume GCN as encoding functions, with its own set of learnable parameters $\Theta$. But, in the experiments, we have also tested fully-connected encoders.
Input :   Graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$
 Adjaceny matrix $\mathbf{A}\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{V}|}$
 Embedding size $K$, Context window $T$,
 Walks length $L$, Regularization $\lambda_{\text{ent}}$
Output :  Embedding matrix $\mathbf{H}\in\mathbb{R}^{|\mathcal{V}|\times K}$
1
21exInit. encoder network $Enc_{\Theta}(*)$;
3 Init. identity matrix features $\mathbf{X}$;
4
5while not converged do
6  Encoding step: $\mathbf{H}\leftarrow\rho(\mathbf{W}^{\top}Enc_{\Theta}(\mathbf{A},\mathbf{X}))$;
7  Sample batch of nodes: $\mathcal{B}\leftarrow Sample(\mathcal{V})$;
8  Init. random walks $\mathcal{W}\leftarrow\emptyset$;
9  foreach $v\in\mathcal{B}$ do
10     Sample random walk: $\mathcal{W}\leftarrow\mathcal{W}\cup RandomWalk(\mathbf{A},v,L)$;
11
12 Random-walk loss: $\mathcal{L}_{\text{rw}}(\mathbf{H},\mathcal{W},T)$;
13  foreach $d\in\{1\dots K\}$ do
14     Aggregate rows of $\mathbf{H}$: $\mathbf{f}_{d}\leftarrow\sum_{v}\mathbf{H}_{vd}$;
15     Compute 1-norm: $|\mathbf{f}|_{1}\leftarrow\sum_{v,d}\mathbf{H}_{vd}$;
16
17 Node affiliation matrix: $\mathbf{F}\leftarrow\mathbf{H}\odot\mathbf{f}$;
18  Disentanglement loss $\mathcal{L}_{\text{dis}}(\mathbf{F})$
19  Regularization loss $\mathcal{L}_{\text{ent}}(\mathbf{F})$
20  Total loss: $\mathcal{L}\leftarrow\mathcal{L}_{\text{rw}}+\mathcal{L}_{\text{dis}}+\lambda_{\text{ent}}\mathcal{L}_{\text{ent}}$;
21  Backpropagate and update $\Theta,\mathbf{W}$;
22
23return $\mathbf{H}$;
[…]
Table A2: Time and space complexity for various embedding methods in terms of number of nodes $|\mathcal{V}|$, number of edges $|\mathcal{E}|$ and latent dimensions $K$. For models trained with random walk sequences, $T$ denotes the context window size and $L$ the random walks length. For models trained with neighborhood sampling, $r$ denotes the number of sampled neighbors per node. For simplicity, in GNNs we consider single-layer architectures and we omit the $\mathcal{O}(K^{2})$ memory usage of model weights since they are negligible compared to storing the embeddings.
Method Time Complexity Space Complexity
DeepWalk $\mathcal{O}(|\mathcal{V}|KTL)$ $\mathcal{O}(|\mathcal{V}|K+|\mathcal{V}|L)$
InfWalk $\mathcal{O}(|\mathcal{V}|^{2}K)$ $\mathcal{O}(|\mathcal{V}|^{2})$
GraphAE $\mathcal{O}(|\mathcal{E}|K+|\mathcal{V}|K^{2})$ $\mathcal{O}(|\mathcal{V}|K)$
GraphSAGE $\mathcal{O}(r|\mathcal{V}|K^{2})$ $\mathcal{O}(r|\mathcal{V}|K)$
Dine $\mathcal{O}(|\mathcal{V}|K^{2})$ $\mathcal{O}(K^{2})$
+base model +base model
DiSeNE $\mathcal{O}(|\mathcal{E}|K+|\mathcal{V}|K^{2}$ $\mathcal{O}(|\mathcal{V}|K+|\mathcal{V}|L)$
$|\mathcal{V}|KTL)$
Our method consists of four main steps:
  1. 1.
[…]
  2. 2.
    Random walk sampling and loss calculation has time/space complexity $\mathcal{O}(|\mathcal{V}|KTL)$ and $\mathcal{O}(|\mathcal{V}|L)$ respectively (Rozemberczki et al., 2019), where $T$ is the context window size and $L$ is the random-walk length (we sample 1 random walk per node, fixing as well the number of negative samples to 1 for each positive sample). $RandomWalk$ function sample a first-order random walk starting from source node $v$ of length $L$.
  3. 3.
    Node affiliation matrix involves computing the entries $\mathbf{F}_{ud}=\sum_{v\in\mathcal{V}_{d}}\phi_{d}(u,v;\mathbf{h})$ as $\mathbf{F}_{ud}=\sum_{v}\mathbf{H}_{ud}\mathbf{H}_{vd}=\mathbf{H}_{ud}\mathbf{f}_{d}$, i.e. by multiplying node embedding entries $\mathbf{H}_{ud}$ with quantities $\mathbf{f}_{d}=\sum_{v}\mathbf{H}_{vd}$. This step involves $\mathcal{O}(|\mathcal{V}|K)$ operations for computing and storing matrix $\mathbf{F}$.
  4. 4.
    Disentanglement and regularization losses involve respectively $\mathcal{O}(|\mathcal{V}|K^{2})$ and $\mathcal{O}(K)$ operations for cosine similarity (matrix products) and entropy (vector sum).
Overall, given that $\mathbf{A}||_{0}$ is $2|\mathcal{E}|$, DiSeNE results in $\mathcal{O}(|\mathcal{E}|K+|\mathcal{V}|K^{2}+|\mathcal{V}|KTL)$ time complexity and $\mathcal{O}(|\mathcal{V}|K+|\mathcal{V}|L)$ space complexity. In Table A2 we compare these findings with computational complexities of competitor methods as they are reported from previous works (Tsitsulin et al., 2021; Chiang et al., 2019; Piaggesi et al., 2024), showing that our approach is in line with established node embeddings (runtime scales linearly with the number of nodes and edges).
Appendix C Downstream Tasks Results for Real Datasets
We tested link prediction for the datasets reported in the main paper. For node classification, we tested PPI and other benchmark biological datasets in multi-label setting (Zhao et al., 2023a): the PCG dataset for the protein phenotype prediction, the HumLoc, and EukLoc datasets for the human and eukaryote protein subcellular location prediction tasks, respectively. Characteristics of additional biological datasets are reported in Table A3. We concatenated node attributes to node embeddings to get an enriched set of predictors that, given our method extract interpretable features, can be used in combination with feature-based explainers (e.g., SHAP) for building fully transparent prediction pipelines. In Figure A1 we report AUC-PR scores for link prediction and node classification in real-world graph data. Generally, scores increase with the number of latent embedding dimensions. Tables A5 and A5 show the maximum scores for link prediction and node classification, demonstrating that our approach can consistently achieve reasonable performances within the expected range of the performance-interpretability trade-off.
@@ Appendix C Downstream Tasks Results for Real Datasets
Figure A1: Downstream tasks results on real-world datasets (link prediction on the top panel, multi-label node classification on the bottom panel) with varying feature dimensions size.
@@
[…]
@@
[…]
Appendix D Explanations for Real Datasets
Figures A2 and  A3 display results for Topological Alignment, Sparsity, Overlap Consistency and Positional Coherence, for real-world datasets. These curves complement the results presented in Tables 1 and 2, by covering a broader range of output embedding sizes $K$ (KK=128 in the main paper) and tuning for the hidden size $D$, as detailed in Section A for Dine (i.e., $D\in[8,16,32,64,128,256,512]$). Generally speaking, DiSe-GAE yields the strongest interpretability scores when the embedding dimension is large, whereas performance differences are less systematic in the low-dimensional regime. Our subsequent analysis therefore focuses on the trade-off between these interpretability metrics and downstream task accuracy (e.g., link prediction).
Figure A4 plots link-prediction AUC-PR against each interpretability metric for all the considered embedding methods. Every cluster of points represents a single method with its range of hidden and output dimensions considered as hyper-parameters. To enable a principled assessment of which method offers the most favorable balance between predictive power and transparency, the left panel of Figure A5 highlights the Pareto frontier in each scatter plot. Models on this frontier are optimal in the sense that no alternative model attains both higher accuracy and greater interpretability, therefore making any other point lying below the frontier strictly dominated with respect to the two criteria, and thus sub-optimal (dominance was assessed while incorporating the statistical uncertainty of each metric, estimated with the standard error from five independent runs per model).
Once the Pareto frontier is identified, we aim to quantify which embedding technique most consistently reaches that frontier. For every method $E$ we collect all of its hyperparameter instantiations $E=\{E_{1},E_{2},\dots E_{m}\}$ (e.g., GraphAEK=4). Let $P$ be the set of all configurations that lie on the frontier. We rank methods with two complementary scores:
$cov(E)=\frac{|P\cap E|}{|P|},\quad\quad\quad eff(E)=\frac{|P\cap E|}{|E|}.$
Coverage is the proportion of methods in the Pareto-frontier that belongs to $E$. Efficiency is the proportion of $E$’s configurations that are Pareto-optimal. Scores reported in Table LABEL:tab:pareto_real combine Coverage and Efficiency via their harmonic mean to have a compact measure quantifying the "average" Pareto-optimality of every embedding method. Using this composite ranking, DiSe-GAE consistently finishes within the top two methods across all datasets.
@@ Appendix D Explanations for Real Datasets
Figure A2: Topological alignment and sparsity results on real datasets with varying feature dimensions size.
Refer to caption Refer to caption
Figure A3: Overlap consistency and positional coherence results on real datasets with varying feature dimensions size.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure A4: Scatter plots with link prediction performance (x-axis) and interpretability metrics (y-axis) on real datasets with varying feature dimensions size.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Table A6: Pareto-optimal methods for real datasets. Dataset Metric #1 #2 Cora Top.Align. DiSe-GAE(.609) GAE+DINE(.130) Sparsity DiSe-GAE(.541) DiSe-FCAE(.165) Over.Cons. DiSe-GAE(.667) DGLFRM(.190) Pos.Coher. DiSe-GAE(.468) GAE+DINE(.213) Wiki Top.Align. DiSe-GAE(.645) GAE+DINE(.194) Sparsity DiSe-GAE(.629) DiSe-FCAE(.135) Over.Cons. DiSe-GAE(.606) DGLFRM(.141) Pos.Coher. DiSe-GAE(.519) GAE+DINE(.198) FB Top.Align. DiSe-GAE(.686) DiSe-FCAE(.235) Sparsity DiSe-GAE(.545) DGLFRM(.156) Over.Cons. DiSe-GAE(.689) GAE+DINE(.067) Pos.Coher. DiSe-GAE(.695) DiSe-FCAE(.168) PPI Top.Align. DiSe-GAE(.505) GAE+DINE(.316) Sparsity DGLFRM(.324) DiSe-GAE(.206) Over.Cons. DiSe-GAE(.769) DGLFRM(.066) Pos.Coher. DiSe-GAE(.469) DGLFRM(.173)
Figure A5: Pareto frontiers for the optimal models considering the interpretability-accuracy trade-offs defined by the four proposed metrics and the link prediction accuracy.
Appendix E Explanations Visualization for Synthetic Datasets
Global Subgraph Explanations
Input :  Graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$
 Embedding function $\mathbf{z}:\mathcal{V}\rightarrow\mathbb{R}^{K}$
 Dimension to explain $d\in\{1\dots K\}$
Output :  Graph mask $\mathbf{M}^{(d)}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$
3
41exInit. graph mask: $\mathbf{M}^{(d)}\leftarrow 0^{|\mathcal{V}|\times|\mathcal{V}|}$;
5 Compute background average attribution: $\zeta_{d}=\frac{1}{|\mathcal{E}|}\sum_{(u,v)\in\mathcal{E}}z_{d}(u)z_{d}(v)$;
6 for $(u,v)\in\mathcal{E}$ do
7  Compute edge attribution: $\phi_{d}(u,v;\mathbf{z})=z_{d}(u)z_{d}(v)-\zeta_{d}$;
8  Add explanation:
9  $\mathbf{M}^{(d)}_{uv}\leftarrow\max\{0,\phi_{d}(u,v;\mathbf{z})\}$;
10
11return $\mathbf{M}^{(d)}$;
[…]
Algorithm 2
[…]
In Figure A6 we show subgraph-level global explanations on synthetic dataset BA-Cliques. Subgraphs are generated for each feature dimension using the procedure described in Section 3.1 (summarized on the left in Algorithm 2) and are based on various unsupervised embedding methods. The explanatory subgraphs demonstrate that our method effectively aligns embedding dimensions with meaningful, non-random functional components of the graph. In contrast, standard methods such as DeepWalk and GraphAE struggle to isolate individual structural units within dimensions. Instead, their embeddings often associate dimensions with groups of cliques or subgraphs that include elements from the random Barabási-Albert scaffold. Additionally, the visualization on the right shows the correlation between latent features, further underscoring that the alignment between embedding dimensions and graph structure is closely tied to the ability to disentangle feature correlation through non-collinearity. Specifcially, we highlight that for DeepWalk and GraphAE the subgraphs exhibit significant overlap, which can be attributed to non-zero correlations within their latent features. In contrast, the uncorrelated features of DiSeNE produce distinct, non-overlapping explanations.
[…]
Figure A6: Subgraph-level global explanations for a representative subset of embedding dimensions, along with corresponding pairwise feature correlation plots, on synthetic dataset BA-Cliques.
Local Subgraph Explanations
Input :  Graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$
 Feature-base explanation matrix $\mathbf{\Psi}\in\mathbb{R}^{|\mathcal{V}|\times K}$
 Dimension to explain $j\in\{1\dots K\}$
Output :  Node mask $\mathbf{B}^{(j)}\in\mathbb{R}^{|\mathcal{V}|}$
3
41exInit. node mask: $\mathbf{B}^{(j)}\leftarrow 0^{|\mathcal{V}|}$;
5 for $v\in\mathcal{V}$ do
6  Add explanation:
7  $B^{(j)}_{v}\leftarrow\max\{0,\Psi_{j}(v;b)\}$;
8
9return $\mathbf{B}^{(j)}$;
[…]
Algorithm 3
[…]
$\text{{(node classification)}}~~~~~b(v)=\sigma(\sum_{j=1}^{K}\beta_{j}h_{j}(v)+\beta_{0})$
$\text{{(link classification)}}~~~~~b(u,v)=\sigma(\sum_{j=1}^{K}\beta_{j}h_{j}(u,v)+\beta_{0})$
Given a vector representation of a graph instance (e.g., a node embedding $\mathbf{h}(v)$ or an edge embedding $\mathbf{h}(u,v)$), and the corresponding prediction from classifier $b$, we compute feature importance with SHAP $\{\Psi^{(\mathcal{V})}_{j}(v;b)\}_{j=1...K}$ or $\{\Psi^{(\mathcal{E})}_{j}(u,v;b)\}_{j=1...K}$ and the corresponding task-based graph masks (we illustrate the pesudo-code for node classification masks in Algorithm 3):
$\mathbf{B}^{(j)}(\mathbf{\Psi}^{(\mathcal{V})})\in\mathbb{R}^{|\mathcal{V}|};~~~B^{(j)}_{v}=\mathrm{max}\{0,\Psi^{(\mathcal{V})}_{j}(v;b)\}$
$\mathbf{B}^{(j)}(\mathbf{\Psi}^{(\mathcal{E})})\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|};~~~B^{(j)}_{uv}=\mathrm{max}\{0,\Psi^{(\mathcal{E})}_{j}(u,v;b)\}$
It is valuable to remark that, training with logistic regression and applying SHAP, the resulting importance scores are simply the coefficients of the regression (Lundberg & Lee, 2017) $\Psi_{j}(x;b)=\beta_{j}(h_{j}(x)-E[h_{j}])$. Thus, combining this methodology to interpretable graph features of DiSeNE, we obtain a fully transparent node/edge classification pipeline for graph data.
Figure A7 present examples of local explanations for node classification tasks on the small-sized synthetic datasets Ba-Cliques and Tree-Grids, using different methods. The experimental settings are consistent with those described in the main paper. On the left, we highlight the local ground-truth structures for the instance nodes depicted in the illustrations. On the right, we display the explanation subgraphs generated by each method, with nodes color-coded according to the respective explanation masks. For GraphAE and DiSeNE, the visualized subgraphs represent the most relevant structures as determined by feature importance attribution from the logistic regression classifier. For GNNExplainer and PGExplainer, the node masks correspond to the algorithm’s output in explaining a 2-layer GCN (Wu et al., 2019). Notably, DiSeNE demonstrates a strong ability to produce meaningful and interpretable node masks, effectively competing with state-of-the-art GNN explanation methods.
[…]
Figure A7: Subgraph local explanations for node classification in BA-Cliques (top) and Tree-Grids (bottom). On the leftmost column, we highlight the local ground-truth structures for the considered instance nodes. On the other columns, we display the explanation subgraphs generated by each method, with nodes color-coded according to the respective explanation masks. For GraphAE and DiSeNE, the visualized subgraphs represent the most relevant structures extracted with Algorithm 3 and determined by feature importance attribution from the logistic regression classifier. For GNNExplainer and PGExplainer, the node masks correspond to the algorithm’s output in explaining a GCN (Wu et al., 2019) in node classification.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure A8: Interpretability metrics computed in Wiki with varying entropy regularization parameters, number of convolutional layers and hidden dimension sizes in DiSe-GAE.
Appendix F Explanations for Synthetic Datasets
Global Explanations
In Figure A9 we plot results for Topological Alignment and Sparsity, on the top and the bottom respectively, on synthetic datasets. Generally, DiSe-FCAE outperforms DiSe-GAE and the other competitors in all the datasets. In Figure A10 we plot results for Overlap Consistency and Positional Coherence, on the top and the bottom respectively, on synthetic datasets. For the overlap metric, DiSe-FCAE and DiSe-GAE consistently outperform the competitors, especially with more than 8 dimensions where they achieve almost perfect overlap. For the positional metric, the competitors GAE+Dine and DW+Dine slightly outperform DiSe methods, especially in large dimensions, while DeepWalk also show good results.
Local Explanations
In Figure A11 we plot results for the plausibility metric on link prediction and node classification, on the top and the bottom respectively, while comparing different unsupervised methods that output node embeddings. Plausibility seems to benefit larger dimension values for DiSe methods and DW+Dine for link prediction. Figure A12 shows the corresponding downstream task accuracy results.
Refer to caption Refer to caption
Figure A9: Topological alignment and sparsity results on synthetic datasets with varying feature dimensions size.